OpenStudy (anonymous):

what does the graph of f'(x) look like if f(x) is the sqrt(6-x)?

OpenStudy (anonymous):

\[f^\prime(x)=-\frac{1}{2}\left(6-x\right)^{-1/2}\]

OpenStudy (anonymous):

I thought it would be 1/2x^-3/2

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

\[f(x)=\sqrt{6-x}\] \[f'(x)=-\frac{1}{2\sqrt{6-x}}\]